A man can row 20km downstream in the time he takes to row 8km upstream. He rows downstream for 1 hr. and 30 min.,then turns and row back for 3 hrs, but he find that he still 3 km from his starting place. Find the rate of the manin still water?
Answer by Paul(988)

20(t-1.50)=8(t+3)
20t-30=8t+24
12t=54
t=4.5
Hence the rate of the man in still water is 4.5km/h